A man is three times as old as his son. Four years ago,he was four times as old as his son was at the time.How old is the son?
Answer by ultrakimidis1(8)

Man's age=3x
Son's age=x
Four years ago,
Man's age =3x-4
Son's age=x-4
Man is 4 times son's age:
3x-4=3(x-4)
Solving this linearly
X=12
Therefore son is 12 years